OpenStudy (anonymous):

express a single logarithm: log c 10+log c 4

OpenStudy (amistre64):

logs add to multiplication: logc(10) + logc(4) = logc(10*4) = logc(40)

OpenStudy (anonymous):

The logarithm of a product is the sum of the logarithms of the factors, hence \[\log_b(xy) = \log_b (x) + \log_b (y). \] So \[\log_c (10) + \log_c (4)=\log_c(4*10) = \log_c(40)\]

OpenStudy (anonymous):

The logarithm of a product is the sum of the logarithms of the factors, hence \[\log_b(xy) = \log_b (x) + \log_b (y). \] So \[\log_c (10) + \log_c (4)=\log_c(4*10) = \log_c(40)\]